OpenStudy (anonymous):

Please help! If the equilibrium constant for a reaction is 0.00010, what does this tell us about the position of equilibrium for that reaction? A. There are more reactants than products at equilibrium. B. There are more products than reactants at equilibrium. C. At equilibrium, the concentration of reactants is the same as that of the products. D. The concentration of the product at equilibrium is 0.00010 M.

OpenStudy (anonymous):

here it shows that kc is moderate hence the reaction occurs in both directions or if you know the initial concentration of product and reactants then there are three cases \[\frac{ product }{ reactant }>kc \] then reaction is backward more reactants are formed \[\frac{ product }{ reactant }<kc \] then reaction is forward more products are formed and \[\frac{ product }{ reactant }=kc \] equilibrium state